# Enhance-A-Video: Better Generated Video for Free

### v1

@@ Abstract
1National University of Singapore  2The University of Hong Kong
3Shanghai Artificial Intelligence Laboratory  4University of Texas at Austin
Figure 1: Enhance-A-Video boosts diffusion transformers-based video generation quality at minimal cost - no training needed, no extra learnable parameters, no memory overhead. Detailed captions are available in Appendix F.
@@ 1 Introduction
Video generation enhancement (He et al., 2024) is designed for addressing the above limitations, where two objectives are primarily considered: ($i$) maintaining temporal consistency across frames, which ensures smooth and coherent transitions, and ($ii$) improving spatial details, which enhances the visual quality of each frame. In UNet-based video generation, VEnhancer (He et al., 2024) designed a video ControlNet (Zhang et al., 2023) to enhance spatial and temporal resolution simultaneously. Nevertheless, the exploration of enhancing DiT-based video generation remains limited, particularly in addressing challenges of temporal consistency and spatial detail preservation.
@@ 3.4 Enhance Block
Finally, the output of the Enhance Block ($\text{CFI}_{\text{enhanced}}$) is utilized to enhance the original temporal attention block output $\mathbf{O}_{\text{attn}}$ in the residual connection (He et al., 2015):
@@ 4.4 Quantitative Analysis
Moreover, we conducted independent evaluations using VBench (Huang et al., 2024) for each video generation model with 5 random seeds. Table 1 shows that integrating Enhance-A-Video consistently improves VBench scores across all models  [^2] VBench may not fully reflect the quality advancements of Enhance-A-Video as discussed in Appendix E.. These results confirm that Enhance-A-Video effectively boosts temporal consistency and visual quality with minimal overhead.
@@ Appendix D More Experimental Results
Appendix E Discussion on VBench
@@
Appendix F Captions for Figure 1

### v2

@@ Abstract
1National University of Singapore  2Shanghai Artificial Intelligence Laboratory  3University of Texas at Austin
Figure 1: Enhance-A-Video boosts diffusion transformers-based video generation quality at minimal cost - no training needed, no extra learnable parameters, no memory overhead. Detailed captions are available in Appendix G.
@@ 1 Introduction
Video generation enhancement (He et al., 2024) is designed for addressing the above limitations, where two objectives are primarily considered: ($i$) maintaining temporal consistency across frames, which ensures smooth and coherent transitions, and ($ii$) improving spatial details, which enhances the visual quality of each frame. In UNet-based video generation (Zhang et al., 2023a; Guo et al., 2024; Xu et al., 2024; Li et al., 2025), Upscale-A-Video (Zhou et al., 2024) integrated a local-global temporal strategy for better temporal coherence, and VEnhancer (He et al., 2024) designed a video ControlNet (Zhang et al., 2023b) to enhance spatial and temporal resolution simultaneously. Nevertheless, the exploration of enhancing DiT-based video generation remains limited, particularly in addressing challenges of temporal consistency and spatial detail preservation.
@@ 3.4 Enhance Block
Finally, the output of the Enhance Block ($\text{CFI}_{\text{enhanced}}$) is utilized to enhance the original temporal attention block output $\mathbf{O}_{\text{attn}}$ in the residual connection (He et al., 2015; Si et al., 2024):
@@ 4.4 Quantitative Analysis
Moreover, we conducted independent evaluations using VBench (Huang et al., 2024) for each video generation model with 5 random seeds. Table 1 shows that integrating Enhance-A-Video consistently improves VBench scores across all models  [^2] VBench may not fully reflect the quality advancements of Enhance-A-Video as discussed in Appendix F.. These results confirm that Enhance-A-Video effectively boosts temporal consistency and visual quality with minimal overhead.
@@ Appendix D More Experimental Results
Appendix E Limitations
Appendix F Discussion on VBench
@@
Appendix G Captions for Figure 1